OpenStudy (anonymous):

Use the elimination method to solve the system. -4x+4y=-8 x-4y=-7

OpenStudy (bahrom7893):

ok let's eliminate x via the 2nd equation: x = -7+4y Plug that into first: -4(-7+4y)+4y=-8

OpenStudy (bahrom7893):

Solve for y: 28-16y+4y=-8 -12y=-36 y = 3

OpenStudy (bahrom7893):

Use that to solve for x

OpenStudy (anonymous):

kk